A person walks 100m in 5 minutes, then 200m in 9 minutes and finally 50



in 4 minutes. Find its average speed?


1
Expert's answer
2022-04-27T16:30:20-0400
"v_{ave}=\dfrac{100m+200m+50m}{5min+9min+4min}=\dfrac{175}{9}m/min"